OpenStudy (anonymous):

M=2/3log 10^21.6/ 10^118? HELP solve for m

OpenStudy (anonymous):

for the same reason that ln(e^x) = x, log(10^x) = x also remember that: log(a/b) = log(a) - log(b) this problem is easier to look at without a calculator if you separate the log quotient into a difference

OpenStudy (anonymous):

assuming its log(10^21.6/10^110) (with brackets)

OpenStudy (anonymous):

answer is 6.5(:

OpenStudy (anonymous):

in solution manual or what you got? C:

OpenStudy (anonymous):

no what i guessed lol